OpenStudy (anonymous):

The line y=3x/4 - 9 is tangent to a circle at (12, 0). The center of the circle is on the line y= 4x/3 - 8. Determine the equation of this circle.

OpenStudy (dumbcow):

equation is: \[(x-9)^{2} + (y-4)^{2} = 25\] you are given slope of tangent line of 3/4 the derivative dy/dx will also be 3/4 at point (12,0) using implicit differentiation \[\rightarrow 2(x-h) +2(y-k) \frac{dy}{dx} = 0\] \[\rightarrow \frac{dy}{dx} = \frac{-(x-h)}{y-k}\] plug in point and set equal to 3/4 ....solve for k in terms of h \[k = 16 - \frac{4}{3}h\] Now we are given line that goes through center (h,k) \[k = \frac{4}{3}h - 8\] thus by substitution: h = 9, k = 4 radius is distance between points (12,0) and (9,4) r = 5

OpenStudy (anonymous):

Fantastic! Thank you so much!

OpenStudy (dumbcow):

your welcome

OpenStudy (zepp):

The tangent of a circle is perpendicular to its radius, so you can find the equation of the radius by using the slope of the tangent.

OpenStudy (zepp):

And then find the distance between where the tangent line crosses the radius line and the center of this circle, you'll get the radius.

OpenStudy (anonymous):

@Montreal_BOY94, c0ckssio, where did you get this question from? Here is an easier understandable solution:  If I draw a line 'c' that is perpendicular to the line 'b', and goes through the point (12, 0), I get the equation of the line 'c' as: y = -3x/4+16 This line also intercept with line 'a' at the center of the circle, thus you find the center as: -3x/4+16 = 4x/3-8 x=9, y=4 DADA